OpenStudy (anonymous):

What are the exact solutions of x2 − 3x − 5 = 0???

OpenStudy (anonymous):

is this the answer "quantity of negative 3 plus or minus the square root of 29 all over 2"

OpenStudy (anonymous):

Try to obtain a square of some expression with x: x^2-3x-5=x^2-3x+9/4-9/4-5=(x-3/2)^2-29/5=0.